Objects, Software Engineering,

and Distributed Systems: A Model and Means for Coexistence

Dr. Charles K. Shank

Department of Computer Engineering, RIT

Abstract

Developing software for distributed architectures with common languages and tools is a challenging task. The developer must be concerned with a wide variety of issues including, concurrency, communications, and resource management. In this talk we present a software development and execution environment, Diamonds, which shows promise in easing the task of creating and running software for such systems.

Diamonds starts from a sound software engineering model that leaves physical design issues cleanly separated from functional ones. This allows the developer to concentrate on application development rather than distributed programming. The mapping of applications to resources in the distributed system is done automatically, in a flexible manner that allows for adaptation to changing resources.

Diamonds is based on a fine-grained, active object model, and includes language, compiler and operating system features which allow the complexity of the underlying distributed system to remain transparent to the user.

Refreshments and cookies will be served